OpenStudy (anonymous):

A point charge of +1.0 µC is moved in the direction of an electric field, and it has a change in electric potential difference energy of 10.0 J. What was the change in electric potential difference? 1 µC = 10^-6 C 1.0 × 108 V 1.0 × 107 V +1.0 × 107 V +1.0 × 108 V

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

The work done moving a charge in an electrical field is given by:\[W=q \Delta V\]where W is work (joules); q is the charge (Coulombs); and ∆V is the change in potential (volts).

OpenStudy (anonymous):

Do you see, now, how to solve the problem?

OpenStudy (anonymous):

What about this question In the absence of friction, gravity, and all other external forces, what kind of speed will an object display if it travels from the bottom to the top of an inclined plane?

OpenStudy (anonymous):

So the only force is the force pushing the object up an inclined plane?

OpenStudy (anonymous):

So would the answer be 1 * 10^7?

OpenStudy (anonymous):

@PsiSquared

OpenStudy (anonymous):

Yes, the answer is B, 1*10^7V=10MV

OpenStudy (anonymous):

Okay, thank you. what about this question? A voltmeter reads 200.0 V between parallel plates that are 2.0 cm apart. What is the strength of the electric field between the two plates? 1.0 × 10^1 N/C 1.0 × 10^2 N/C 1.0 × 10^3 N/C 1.0 × 10^4 N/C

OpenStudy (anonymous):

@psisquared

OpenStudy (anonymous):

@IhteshamMalik

OpenStudy (anonymous):

\[E = - \frac{ \Delta v }{ \Delta r }\]

OpenStudy (anonymous):

E = 200/ 0.02 = \[1 \times 10^{3} \]

OpenStudy (anonymous):

SO the answer is third one. and Volt / meter = Newton/coulomb

OpenStudy (anonymous):

10,000 N per coulomb.

OpenStudy (anonymous):

Oh, okay! that makes sense.

OpenStudy (anonymous):

sense always makes sense... :D

OpenStudy (anonymous):

lol, :) thanks for the help!

OpenStudy (anonymous):

You're welcome bro.

OpenStudy (anonymous):

what about this question? If 4.0 × 1018 J of work done on a point charge changes its electric potential difference by 25.0 V, then what is the magnitude of the charge? 1.6 × 101^9 C 4.0 × 101^9 C 1.6 × 101^8 C 4.0 × 101^8 C

OpenStudy (anonymous):

@IhteshamMalik

OpenStudy (anonymous):

SO you are not gonna to sleep??? :D

OpenStudy (anonymous):

\[q = \frac{ W }{ \Delta V }\]

OpenStudy (anonymous):

@lmckenzie3

OpenStudy (anonymous):

????

OpenStudy (anonymous):

sorry, my interet connection went out for a while. I'm here. No, im not sleeping or going to go to sleep. haha

OpenStudy (anonymous):

@IhteshamMalik

OpenStudy (anonymous):

okay solve your question then with above formula...

OpenStudy (anonymous):

Okay, so would th answer be 1.6 * 10^-19?

OpenStudy (anonymous):

ooo how did you do this???

OpenStudy (anonymous):

Q = W/V = \[q = \frac{ 4 \times10^{18} }{ 25 }\]

OpenStudy (anonymous):

i did exactly like the formula above but came up with the answer 1.6 * 10^-19 on my calculator.

OpenStudy (anonymous):

Did i wrote right value of work done???

OpenStudy (anonymous):

and check your answers also

OpenStudy (anonymous):

The answers i typed above are incorrect. I just now realized that. The choices are supposed to be 1.6 × 10^19 C 4.0 × 10^19 C 1.6 × 10^18 C 4.0 × 10^18 C

OpenStudy (anonymous):

i have answer. \[1.6 \times 10^{17} C\] and i have no doubt about answer.

OpenStudy (anonymous):

Check question also, i think \[W = 4 \times 10^{19}\]

OpenStudy (anonymous):

hmm what you thinking??

OpenStudy (anonymous):

Oh you are right thank you! can you help me with this other question? For the electric field shown, which of the following statements describes an increase in voltage? A positive charge moves along an equipotential line. A positive charge crosses an equipotential line toward the source charge. A negative charge crosses an equipotential line toward the source charge. You cannot tell from the information given.

OpenStudy (anonymous):

I know the answer is either A or B. I just am stuck between the two and dont know which one it is

OpenStudy (anonymous):

Point B is true.

OpenStudy (anonymous):

Because we have the equation \[ V = k \frac{ q }{ r }\]

OpenStudy (anonymous):

So increase in distance from the source, causes decrease in potential.